OpenStudy (anonymous):

you made model of the space needle in seatle washington for a report on architecture in the united state you use a scale of 1 in. 50 ft. your model is 12 in.1 inch tall estimate the actual height of the space needle

OpenStudy (anonymous):

is it supposed to be 12 ft 1 in tall?

OpenStudy (anonymous):

inches tall

OpenStudy (anonymous):

your model is 12 in.1 inch tall does not make sense.

OpenStudy (anonymous):

12.1 inches tall

OpenStudy (anonymous):

the book said that

OpenStudy (anonymous):

it's 8 grade algebra 1

OpenStudy (anonymous):

ok, that makes more since. so 12.1(50) = 12(50) + .1(50) = 600 + 5 = 605 605 feet should be your answer